Given the following information and formulae from the lecture, calculate earned value cost variance, and schedule...

  1. Given the following information and formulae from the lecture, calculate
    • earned value
    • cost variance, and
    • schedule variance:

Rate of performance: 45%

Planned value: $20000

Actual cost: $40000
Answer #1

WE HAVE THE FOLLOWING VALUES:
PV = 20000
RATE OF PERFORMANCE = 45
ACTUAL COST = 40000

FORMULA:
EARNED VALUE = PLANNED VALUE * RATE OF PERFORMANCE
COST OF VARIANCE = EARNED VALUE - ACTUAL COST
SCHEDULE VARIANCE = EARNED VALUE - PLANNED VALUE

EV = 20000 * 0.45 = 9000
CV = 9000 - 40000 = -31000
SV = 9000 - 20000 = -11000